Top Company holds 90 percent of Bottom Company’s common stock. In the current year, Top reports sales of $886,000 and cost of goods sold of $664,500. For this same period, Bottom has sales of $368,000 and cost of goods sold of $220,800. During the current year, Bottom sold merchandise to Top for $195,000. The parent still possesses 35 percent of this inventory at the current year-end. Bottom had established the transfer price based on its normal gross profit rate. What are the consolidated sales and cost of goods sold?

  • $1,059,000 and $758,550.

  • $1,059,000 and $717,600.

  • $1,059,000 and $713,050.

  • $1,022,200 and $912,600.
Answer #1

$1,059,000 and $717,600 is the correct option.

Reason:-

Consolidated sales :-

= 886,000 + 368,000 - 195,000 = $1,059,000

Consolidated Cost of goods sold :-

Bottom gross profit ratio = (368,000 - 220,800)/368,000 = 40%

Top's inventory from Bottom = 195,000 x 35% = $68,250

Unrealized gain = 68,250 x 40% = $27,300

Cost of goods sold = 664,500 + 220,800 - (195,000 - 27,300) = $717,600